# Towards properties of adversarial image perturbations

### v1

@@ Introduction
One of the important problems when analyzing the interaction of neural networks with adversarial perturbations is ability to generate these perturbations in the data. The majority of methods for producing adversarial examples rely on gradient-based optimization, however, some use other optimization techniques. Among them the authors of [2] proposed the Fast Gradient Sign Method (FGSM) in which the adversarial perturbation is generated using the cross-entropy as a loss function and $l_{\infty}$ norm for measuring amplitudes of perturbations. The similar method called Fast Gradient Method (FGM) exploits the same loss function in combination with $l_{2}$ norm [5]. In the current work we use the same gradient-based approach and investigate the perturbations applying the same two norms. One important distinction of this work results from the use of different loss function. For this purpose we apply VMAF which, being a trainable machine learning model, also was shown to be vulnerable to adversarial perturbations. Unlike neural networks, in the case of image quality measuring models adversarial perturbations may not only be generated by some gradient-like method, but can also emerge as a result of image processing with already existing popular methods for image quality improvement. For example, in [6] the authors apply various well-known image processing techniques, such as gamma correction, adaptive histogram equalizaion (CLAHE), unsharp masking and show that VMAF score can be significantly improved on average (by $\approx 60\%$). Many of these techniques give the highest gains when the image is substantially corrupted. This provides some evidence that the VMAF model diverges with the human estimate of image quality, indicating potential vulnerability of VMAF to adversarial perturbations.
@@
Thus, the goal of the present paper is twofold: 1) investigate and present properties of perturbations generated on images by means of direct VMAF optimization; 2) provide more evidence for applicability of the algorithm implementing VMAF proposed in [7]. For the latter goal we also consider the problem of image restoration from the point of view previously studied in [9], where various metrics were compared among themselves in terms of the quality of image restoration. Without pretending on noticeably new results at this particular point, we restore images using VMAF-torch numerically and demonstrate that our results are in correspondence with [9]. It may be worth noting at the outset that the question of the loss function properties wrt. adversarial perturbations is not studied in this paper and is assumed to be presented elsewhere: we only focus on the properties of the adversarial perturbations.
@@
[…]
@@
It can be seen than perturbing pixels of the image with $\delta$ by as little as one pixel value can result in VMAF growth up to $10$ units. It is also noticeable that the dependence of the gain on the perturbation amplitude is approximately power-law with the exponent $\approx 2.2$, though we do not have additional evidence for any power-law behavior here. On the other hand, Fig. 1 for different $\varepsilon$ shows that the results on the VMAF growth are hardly dependent on the data set: all three data sets demonstrate very close behavior. The perturbed image $R+\delta$ along with the reference $R$ and the perturbation $\delta$ image are shown in Fig. 2111We need to have a look at deltas for $\epsilon=1,2,4,8,16$ etc. We need to have a look at the spectra in each case. It would be reasonable to evaluate the information loss in each case. It also would be profitable to compare the mutual information with that for gaussian noise. We also see that the perturbation acts as a high-pass filter. This may be studied further..
@@
$=10\log\left(\frac{255^{2}}{\frac{1}{mn}\|R-(R+\delta)\|^{2}_{2}}\right)=10\log\left(\frac{255^{2}mn}{\|\delta\|^{2}_{2}}\right),$
@@
Overall it is seen that the perturbation of the image spectrum magnitude produced by the adversarial perturbation in the pixel domain is insignificant, which is also seen from the amplitude of the spectrum for the perturbation which is at least two orders of magnitude smaller than the reference and perturbed spectra (Fig. 6). The spectrum for the perturbation image may account for the observations in Fig. 4c. It is seen that the tails of the spectra both for the reference and perturbation images exhibit similar behavior, keeping the mutual structure of frequencies. This corresponds to smaller objects, edges, lines etc. observed in Figs. 4, thereby reproducing the objects in the reference image. On the other hand, the low frequency part of the spectrum is noticeably reduced for the perturbation signal, which corresponds to disappearance of larger structures compared to the reference image. Overall, the behavior of the spectra is close to the linear, which is in good correspondence to the classical results concerning the spectra for natural images [12]. At the same time the structure of the spectra shown in Fig. 6 implies the corresponding correlation function incompatible to the independent noise, which provides an additional evidence of the compex nature of perturbations.
Despite these observations the computations show that 1D power spectrum for a sample of images is weakly sensitive to VMAF growth (cf. blue and red curves in Fig. 6). Realistic VMAF variations resulting from standard enhancement methods may provide improvements $\sim 4-6$ VMAF units, while for the perturbed image in Fig. 4 VMAF attains the magnitude $134$, significant increase, taking into account that the normal range for high-quality images is $\sim 85-95$ VMAF units. This significant increase, however, is hardly can be viewed in the spectra in Fig 6.
@@
Figure 7: Average VMAF gain achieved by the unsharp masking filter (size $5\times 5$), CLAHE, and adversarial perturbations using two norms $l_{\infty}$ and $l_{2}$ on Netflix dataset for comparable PSNR values. Different points on the red and blue curves represent different values of amplification parameters for CLAHE and the unsharp masking correspondingly; we set these parameters so that the processed images fall withing PSNR range of $40-50$ dB. The points on the green curve correspond to various PSNR lower bounds established when training the perturbation model. The points on the violet curve correspond to the maximum perturbation magnitudes $\varepsilon$, shown with the numbers near each point.
@@
The discussion concerning the perturbations and the corresponding VMAF growth presented in the previous sections was quite generic. To try to investigate the perturbations in connection with some more practical task of image processing we address the problem of image recovery. We would like to take an attempt to apply VMAF and its submetrics to the problem of image recovery from noise. One of the goals pursued here is a comparison with the results obtained in [9], where a number of metrics were used for the same purpose but in our approach we can directly optimize over VMAF using the implementation of VMAF on Pytorch which we presented recently[7] [^3] The implementation is available opensource at [8].. Note, the the other tasks studied in [9], such as denoising or deblurring and different FR metrics ranking are not considered here; we restrict ourselves with the analysis of image restoration problem using VMAF implementation.
@@
Figure 9: 1D power spectra for the reference and restored images obtained by $VIF_{0}$ and VMAF optimizations and corresponding to Fig. 10 a), d), h) correspondingly. Also 1D power spectra obtained for the images obtained by $VIF_{0}$ and VMAF metrics maximization and corresponding to Fig. 11 d), h) are shown. The spectra were computed by the method explained in the caption to Fig. 6.
@@
For the case of initial image being the compressed version of the reference one, the results are shown in Fig. 12. Now VIF at scale 00 is actually able to restore the details of the image lost due to compression. However, some high-frequency noise is added to the image; this can be caused by the fact that image is smoothed by a gaussian filter, so that noisy and non-noisy version of the image are indistinguishable to VIF. VMAF, ADM and VIF for other scales are not able to recover the image and also add additional artifacts.
@@ Discussion
One of the natural questions which can be raised in connection with this work is the problem of so called metric breaking. Simplistically speaking, it poses the following problem. If we have a processing method which alters the image, and if we use VMAF or any other quality measure as our main means to judge about the comparable quality of the image before and after the alteration, it is claimed that one can “deceive” the metric trying to improve VMAF score by applying various techniques, e.g., those described above. Putting aside the jargon (metric breaking, deceive etc.), such problem formulation seems to be too narrow and unilateral. Objectively, we rely our judgements about the image quality on some model which computes some score for us. We then can compare this score to our subjective judgement on whether the image is good or bad [^6] We do not take into account more sophisticated methods for comparison subjective and objective judgements such as mean opinion score (MOS).. So, first of all, it seems reasonable to stress that any (parametric) model is only restrictedly applicable even in the sphere it was designed for. Probably, as it often happens, in some parameter regime the model declines from the phenomenon it endevours to model, and we would inevitably obtian irrelevant results. It does not mean, however, that the model is bad; it merely means it has some sphere of application and has to be used cautiously.
@@
But it seems the problem we tackle dealing with adversarial examples in the field of machine learning is subtler. One argues that using some small deviation from the observed data the model behaves inadequately (provides wrong classification, abnormally high score etc.). This is often expressed by saying that the model is not robust. This is clear for the classificastion tasks where this can imply that some small perturbation of data changes the class to which the data is assigned. However, the problem we deal with is not that of classification. Instead, we observe gradual changes in VMAF score when increasing the perturbation amplitude. First of all, the natural question here is what the smallness of deviations mean. Specifically, we have to decide in what norm we measure the smallness. Even if the norm is specified (there is a tendency to rely on some eucleadin norm, like $l_{\infty}$ or $l_{2}$ as a measure of closeness) one of the problems when dealing with adversarial examples is that image perturbations may not be perceivable with the human eye. In other words, an image can be judged as good subjectively, but after some small changes in terms of the specified norm it is still evaluated as close to the previous one, while its objective score changes significantly. This only may mean that the subjective judgement of what is small and what is big deviation or, to put it another way, the putative subjective norm underlying our judgement is not quite suitable for the model, i.e., is not in good correspondence with typically used norms and that the sensitivity or the resolution power of human visual system (HVS) is not aligned to the model in question. Thus, it seems the question of seeking more suitable norms still remains.
@@
There is also another point to consider. The data presented to the model is implied to be generated by some natural process. We mean, for example, the image our retinal cells normally obtain is constructed as some natural process in the external environment [^7] Various forms of animation, though presenting synthetic images, are viewed here as natural-like ones.. The question then is whether some deviations or perturbations, even small ones in some sense, keep us close to the data likely occurring in nature (Cf. [2], p. 2). If this is not the case, then the divergence of the model and subjective judgement may be accounted for by the same arguments as provided above. If, on the other hand, naturally occurring data do result in inadequate model behavior, the problem becomes more complicated. But whether we have evidence of naturally occurring images extemely close to the other ones subjectively and providing noticeably different quality scores?
@@ Conclusions
In this work we provided the analysis of adversarial perturbations based on a differentiable version of VMAF implemented in the Pytorch framework. Using this implementation we tried to investigate statistical properties of the adversarial perturbations generated by the stochastic gradient algorithm using VMAF as a loss function in terms of $l_{2}$ and $l_{\infty}$ norms. Among several observations made in this work, those worth reiterating here are of weak linear dependence of the perturbations amplitudes on the brightness in the pixels. This observation can be accounted for by edge sharpening effects ususally found when increasing VMAF. On the other hand, this observation can be viewed as a zero-order approximating model of adversarial perturbations. Being purely empirical, this model requires further analysis as a potential means for perturbations generation. We presume to elaborate this further.

### v2

@@ Introduction
One of the important problems when analyzing the interaction of neural networks with adversarial perturbations is ability to generate these perturbations in the data. The majority of methods for producing adversarial examples rely on gradient-based optimization, however, some use other optimization techniques. Among them the authors of [2] proposed the Fast Gradient Sign Method (FGSM) in which the adversarial perturbation is generated using the cross-entropy as a loss function and $l_{\infty}$ norm for measuring amplitudes of perturbations. The similar method called Fast Gradient Method (FGM) exploits the same loss function in combination with $l_{2}$ norm [5]. In the current work we use the same gradient-based approach and investigate perturbations applying the same two norms. One important distinction of this work results from the use of different loss function. For this purpose we apply VMAF which, being a trainable machine learning model, also was shown to be vulnerable to adversarial perturbations. Unlike neural networks, in the case of image quality measuring models adversarial perturbations may not only be generated by some gradient-like method, but can also emerge as a result of image processing with already existing popular methods for image quality improvement. For example, in [6] the authors apply various well-known image processing techniques, such as gamma correction, adaptive histogram equalizaion (CLAHE), unsharp masking and show that VMAF score can be significantly improved on average (by $\approx 60\%$). Many of these techniques give the highest gains when the image is substantially corrupted. This provides some evidence that the VMAF model diverges with the human estimate of image quality, indicating potential vulnerability of VMAF to adversarial perturbations.
@@
Thus, the goal of the present paper is twofold: 1) investigate and present properties of perturbations generated on images by means of direct VMAF optimization; 2) provide more evidence for applicability of the algorithm implementing VMAF proposed in [7]. For the latter goal we also consider the problem of image restoration from the point of view previously studied in [9], where various metrics were compared among themselves in terms of the quality of image restoration. Without pretending to noticeably new results at this particular point, we restore images using VMAF-torch numerically and demonstrate that our results are in correspondence with [9]. It may be worth noting at the outset that the question of the loss function properties wrt. adversarial perturbations is not studied in this paper and is assumed to be presented elsewhere: we only focus on the properties of the adversarial perturbations.
@@
where $\|\cdot\|$ is some norm on a vector space and $\delta\in\mathbf{R}^{d}$ may be referred to as perturbation. Common choices for the norm in the above expression are $l_{\infty}$ norm and $l_{2}$ norm; we generate perturbations for both. It may be worth stressing from the beginning that in this formulation the problem implies that the signal $R$ is perturbed additively, and thus the problem can be thought of as that of a quality metric maximization by means of an additive component, which, in a sense, plays the role of a noise, in the same way as the task of signal-to-noise ratio (SNR) maximization in the presence of gaussian additive noise can be formulated. Another alternative to this formulation would be application of a filter to the original signal $R$, i.e., a perturbation which is obtained by convolving the original signal with some other signal instead of addition.
@@
It can be seen that perturbing pixels of the image with $\delta$ by as little as one pixel value can result in VMAF growth up to $10$ units. It is also noticeable that the dependence of the gain on the perturbation amplitude is approximately power-law with the exponent $\approx 2.2$, though we do not have additional evidence for any power-law behavior here. On the other hand, Fig. 1 shows that the results on the VMAF growth are hardly dependent on the data set for different $\varepsilon$: all three data sets demonstrate very close behavior. The perturbed image $R+\delta$ along with the reference $R$ and the perturbation $\delta$ image are shown in Fig. 2.
@@
$10\log\left(\frac{255^{2}}{\frac{1}{mn}\|R-(R+\delta)\|^{2}_{2}}\right)=10\log\left(\frac{255^{2}mn}{\|\delta\|^{2}_{2}}\right),$
@@
Overall it is seen that the perturbation of the image spectrum magnitude produced by the adversarial perturbation in the pixel domain is insignificant, which is also seen from the amplitude of the spectrum for the perturbation which is at least two orders of magnitude smaller than the reference and perturbed spectra (Fig. 6). The spectrum for the perturbation image may account for the observations in Fig. 4c. It is seen that the tails of the spectra both for the reference and perturbation images exhibit similar behavior, keeping the mutual structure of frequencies. This corresponds to smaller objects, edges, lines etc. observed in Figs. 4, thereby reproducing the objects in the reference image. On the other hand, the low frequency part of the spectrum is noticeably reduced for the perturbation signal, which corresponds to disappearance of larger structures compared to the reference image. Overall, the behavior of the spectra is close to linear, which is in good correspondence to the classical results concerning the spectra for natural images [12]. At the same time the structure of the spectra shown in Fig. 6 implies the corresponding correlation function incompatible to the independent noise, which provides an additional evidence of the complex nature of the perturbations.
Despite these observations the computations also show that 1D power spectrum for a sample of images is weakly sensitive to VMAF growth (cf. blue and red curves in Fig. 6). Realistic VMAF variations resulting from standard enhancement methods may provide improvements $\sim 4-6$ VMAF units, while for the perturbed image in Fig. 4 VMAF attains the magnitude $134$, significant increase, taking into account that the normal range for high-quality images is $\sim 85-95$ VMAF units. This significant increase, however, is hardly can be viewed in the spectra in Fig 6.
@@
Figure 7: Average VMAF gain achieved by the unsharp masking filter (size $5\times 5$), CLAHE, and adversarial perturbations using two norms $l_{\infty}$ and $l_{2}$ on Netflix dataset for comparable PSNR values. Different points on the red and blue curves represent different values of amplification parameters for CLAHE and the unsharp masking correspondingly; we set these parameters so that the processed images fall withing PSNR range of 405040-50 dB. The points on the green curve correspond to various PSNR lower bounds established when training the perturbation model. The points on the violet curve correspond to the maximum perturbation magnitudes $\varepsilon$, shown with the numbers near each point.
@@
The discussion concerning the perturbations and the corresponding VMAF growth presented in the previous sections was quite generic. To try to investigate the perturbations in connection with some more practical task of image processing we address the problem of image recovery from noise. One of the goals pursued here is a comparison with the results obtained in [9], where a number of metrics were used for the same purpose but in our approach we can directly optimize over VMAF using the implementation of VMAF on Pytorch which we presented recently[7] [^2] The implementation is available opensource at [8].. Note that other tasks studied in [9], such as denoising or deblurring and different FR metrics ranking are not considered here; we restrict ourselves with the analysis of image restoration problem using VMAF implementation.
@@
Figure 9: 1D power spectra for the reference and restored images obtained by $VIF_{0}$ and VMAF optimizations and corresponding to Fig. 10 a), d), h) correspondingly. In addition, 1D power spectra obtained for the images obtained by $VIF_{0}$ and VMAF metrics maximization and corresponding to Fig. 11 d), h) are shown. The spectra were computed by the method explained in the caption to Fig. 6.
@@
For the case of initial image being the compressed version of the reference one, the results are shown in Fig. 12. Now VIF at scale 0 is actually able to restore the details of the image lost due to compression. However, some high-frequency noise is added to the image; this can be caused by the fact that image is smoothed by a gaussian filter, so that noisy and non-noisy version of the image are indistinguishable to VIF. VMAF, ADM and VIF for other scales are not able to recover the image and also add additional artifacts.
@@ Discussion
One of the natural questions which can be raised in connection with this work is the problem of so called metric breaking. Simplistically speaking, it poses the following problem. If we have a processing method which alters the image, and if we use VMAF or any other quality measure as our main means to judge about the relative quality of the image before and after the alteration, it is claimed that one can “deceive” the metric trying to improve VMAF score by applying various techniques, e.g., those described above. Putting aside the jargon (metric breaking, deceive etc.), such problem formulation seems to be too narrow and unilateral. Objectively, we rely our judgements about the image quality on some model which computes some score for us. We then can compare this score to our subjective judgement on whether the image is good or bad [^5] We do not take into account more sophisticated methods for comparison subjective and objective judgements such as mean opinion score (MOS).. So, first of all, it seems reasonable to stress that any (parametric) model is only restrictively applicable even in the sphere it was designed for. Probably, as it often happens, in some parameter regime the model declines from the phenomenon it endevours to model, and we would inevitably obtain irrelevant results. It does not mean, however, that the model is bad; it merely means it has some sphere of application and has to be used cautiously.
@@
But it seems the problem we tackle dealing with adversarial examples in the field of machine learning is subtler. One argues that using some small deviation from the observed data the model behaves inadequately (provides wrong classification, abnormally high score etc.). This is often expressed by saying that the model is not robust. This is clear for the classification tasks where this can imply that some small perturbation of data changes the class to which the data is assigned. However, the problem we deal with is not that of classification. Instead, we observe gradual changes in VMAF score when increasing the perturbation amplitude. First of all, the natural question here is what the smallness of deviations mean. Specifically, we have to decide in what norm we measure the smallness. Even if the norm is specified (there is a tendency to rely on some euclidean norm, like $l_{\infty}$ or $l_{2}$ as a measure of closeness) one of the problems when dealing with adversarial examples is that image perturbations may not be perceivable with the human eye. In other words, an image can be judged as good subjectively, but after some small changes in terms of the specified norm it is still evaluated as close to the previous one, while its objective score changes significantly. This only may mean that the subjective judgement of what is small and what is big deviation or, to put it another way, the putative subjective norm underlying our judgement is not quite suitable for the model, i.e., is not in good correspondence with typically used norms and that the sensitivity or the resolution power of human visual system (HVS) is not aligned to the model in question. Thus, it seems the question of seeking more suitable norms still remains.
@@
There is also another point to consider. The data presented to the model is implied to be generated by some natural process. We mean, for example, the image our retinal cells normally obtain is constructed as some natural process in the external environment [^6] Various forms of animation, though presenting synthetic images, are viewed here as natural-like ones.. The question then is whether some deviations or perturbations, even small ones in some sense, keep us close to the data likely occurring in nature (Cf. [2], p. 2). If this is not the case, then the divergence of the model and subjective judgement may be accounted for by the same arguments as provided above. If, on the other hand, naturally occurring data do result in inadequate model behavior, the problem becomes more complicated. But whether we have evidence of naturally occurring images extremely close to the other ones subjectively and providing noticeably different quality scores?
@@ Conclusions
In this work we provided the analysis of adversarial perturbations based on a differentiable version of VMAF implemented in the Pytorch framework. Using this implementation we tried to investigate statistical properties of the adversarial perturbations generated by the stochastic gradient algorithm using VMAF as a loss function in terms of $l_{2}$ and $l_{\infty}$ norms. Among several observations made in this work, those worth reiterating here are of weak linear dependence of the perturbations amplitudes on the brightness in the pixels. This observation can be accounted for by edge sharpening effects usually found when increasing VMAF. For this explanation we provided sufficient evidence. On the other hand, this observation can be viewed as a zero-order approximating model of adversarial perturbations. Being purely empirical, this model requires further analysis as a potential means for perturbations generation. We presume to elaborate this further.
@@
Data availability statement
The authors declare that the data supporting the findings of this study are available through the links provided in the reference list.